# Partial fraction decompositions on hyperplane arrangements

### v1

@@ Abstract.
We initiate the study of partial fraction decompositions (PFDs) in several variables using tools from commutative algebra. We give criteria for when a rational function with poles on a hyperplane arrangement has a desirable PFD. Our criteria are obtained by examining the primary decomposition of ideals coming from hyperplane arrangements. We then present an algorithm for finding a PFD that satisfies properties desired by physicists, and demonstrate the effectiveness of this algorithm for computing large examples coming from Feynman integrals.
@@ 1. Introduction
Partial fraction decompositions (PFDs) of rational functions in several variables appear frequently in physics. Examples include abstract applications, such as the re-casting of scattering amplitudes as canonical forms of polytopes [ABL], and computational applications like developing efficient algorithms to compute Feynman integrals [HvM]. Despite their ubiquity in physics, studies of PFDs seldom appear in mathematical literature [BBD+, Bec, Lei]. There has been interest from physicists to utilize algebro-geometric techniques in order to compute certain PFDs [BW+, CDE, DP]. In this paper we study PFDs from the perspective of commutative algebra, by using primary decomposition and algorithms based upon Gröbner bases.
@@
In applications to physics, one would like to know when a rational function $f/g$ has a PFD, and if so, how to make the PFD “optimal.” Here, optimal can refer to a number of parameters, including minimizing the number of terms $m$ in the PFD, maximizing the degrees of the numerators $f_{i}$, and avoiding spurious poles (the introduction of such poles can complicate calculations and obscure physical properties and symmetries [Hod]).
@@
Let $f\in K[x_{1},\ldots,x_{r}]$ be a homogeneous polynomial of degree $\deg(f)\geq d\geq 1$, and let $\ell_{1},\ldots,\ell_{n}$ be nonzero linear forms with $n\geq d$. Then $f/(\ell_{1}\cdots\ell_{n})$ has a PFD of degree $d$ as in (1) if and only if $f$ vanishes to order at least dn+|S|d-n+|S| on all flats $S$ of size at least nd+1n-d+1 of the hyperplane arrangement associated with the linear forms.
Our approach to proving Theorem 1.2 is with commutative algebra: given linear forms $\ell_{1},\ldots,\ell_{n}$ of the polynomial ring $R=K[x_{1},\ldots,x_{r}]$, we study the primary decomposition of the following ideal:
@@
Existence of a PFD of order $d$ for the rational function $f/(\ell_{1}\cdots\ell_{n})$ is equivalent to the numerator $f$ being in the ideal $I_{L,d}$, and our primary decomposition gives an equivalent geometric condition for when this occurs.
One can use Macaulay2 [M2] to determine the existence of and algorithmically compute PFDs of the form in (1). Our algorithm (Algorithm 2) for finding a PFD is based upon Gröbner basis algorithms. It satisfies conditions (i), (ii), (iii), and (iv) of the “wishlist” for a good partial fraction algorithm, described by Heller and von Manteuffel in [HvM, §2.3] and based upon work of Leĭnartas [Lei]:
@@
We now discuss the organization and contributions of this paper. In §2, we study the ideals $I_{L,d}$ in (2) defined from hyperplane arrangements. We give primary decompositions for these ideals (Corollary 2.9 and Theorem 2.12), which are presented in terms of the matroid associated with the hyperplane arrangement. Our primary decomposition also allows us to see that the ideals $I_{L,d}$ have linear resolutions (Corollary 2.11). We then apply our decomposition in §3 to study the ideals $I_{L,d}$ where $L$ is the braid arrangement. In §4, we also use our primary decomposition to deduce vanishing behavior of adjoint polynomials arising from polytopes; these polynomials are objects of much study in physics, combinatorics, and algebraic geometry. We then return to our main motivation of PFDs in §5: in this section, we give criteria for when PFDs of degree $d$ exist (Theorems 5.2 and 5.3). We present Algorithms 1 and 2 to determine the existence of and compute PFDs; we also demonstrate how our algorithms satisfy all four conditions (i)(iv) of a good PFD algorithm. In §6, we then show how our criteria and algorithm can be applied to PFDs coming from two different contexts from physics: Feynman integrals (§6.1) and flat-space wavefunction coefficients (§6.2). We conclude in §7 with remarks on the outlook and directions for future investigation for the algebraic study of PFDs.
@@ 2. Primary decompositions
Let $d$ be a positive integer with $1\leq d\leq n$. Given a hyperplane arrangement $L=(\ell_{1},\ldots,\ell_{n})$, we study the ideal $I_{L,d}\subset R$ defined by $d$-fold products of the linear forms:
@@
We allow for the cases that some of the $\ell_{i}=0$ are zero, or that the $\ell_{i}$ may not all be distinct, i.e., $\ell_{i}=\ell_{j}$ for for some $i\neq j$. However, we assume that $I_{L,d}\neq 0$ is not the zero ideal, so not all of the $\ell_{i}=0$ are zero. The main results of this section give a primary decomposition for $I_{L,d}$.
2.1. Generic hyperplane arrangements
@@
To state the following proposition, we introduce some notation. Write each linear form $\ell_{i}\in R$ in the form $\ell_{i}=\sum_{j=1}^{r}a_{ij}x_{j}$, with $a_{ij}\in K$. Define $A=(a_{ij})$ to be the $(n\times r)$-matrix of these coefficients. For $1\leq i_{1}<\cdots<i_{r}\leq n$, denote by $A[i_{1},\ldots,i_{r}]$ the $r$-minor of AA obtained by taking the submatrix of AA with rows labeled by $\ell_{i_{1}},\ldots,\ell_{i_{r}}$. Now, define $\mathcal{M}$ to be the $(\binom{n}{d}\times\binom{n}{d})-$matrix of coefficients, with rows labeled by $d$-fold products $\ell_{i_{1}}\cdots\ell_{i_{d}}$ of the linear forms, and columns labeled by degree $d$ monomials of $R$.
@@
The flats of this matroid are the subsets of the ground set of cardinality at most $r$, together with the ground set itself. This foreshadows Corollary 2.9 in the following subsection, where we consider arbitrary hyperplane arrangements.
$2.$ Arbitrary projective case
In this subsection, let $r,n,d$ be positive integers with $n\geq d$. We consider arbitrary hyperplane arrangements determined by $n$ linear forms $L=(\ell_{1},\ldots,\ell_{n})$, with $\ell_{i}\in R=K[x_{1},\ldots,x_{r}]$. Our main result (Corollary 2.9) is a primary decomposition for the ideal $I_{L,d}$.
Our argument follows that of [CH, §3], in which the authors study products of ideals generated by linear forms; we note that although of a similar flavor, our ideal $I_{L,d}$ cannot be written as the product of such ideals.
Definition 2.5.
Let II be a homogeneous ideal in the polynomial ring $R$, and let $\mathfrak{m}=\langle x_{1},\ldots,x_{r}\rangle$ be the homogeneous maximal ideal. We define the saturation $I^{\operatorname{sat}}$ of II to be the ideal:
$I^{\operatorname{sat}}=\{f\in R:f\cdot\mathfrak{m}^{p}\subset I\text{ for some }p\in\mathbb{N}\}.$
The saturation degree $\mathrm{sat}(I)$ of II is then defined to be the smallest integer $q\in\mathbb{N}$ such that for any $f$ homogeneous with $\deg(f)\geq q$:
$f\in I\quad\Leftrightarrow\quad f\in I^{\operatorname{sat}}.$
For $S\subset[n]$, let $I_{S}=\langle\ell_{i}:i\in S\rangle$ be the prime ideal generated by a subset $S$ of the linear forms. Notice that any power of $I_{S}$ is an $I_{S}$-primary ideal; our primary decomposition of $I_{L,d}$ consists of these $I_{S}$-primary ideals.
Lemma 2.6.
We have the following inclusion of ideals:
$I_{L,d}\subset\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}.$
Proof.
Fix $S\subset[n]$ with $|S|\geq n-d+1$. It suffices to show that $I_{L,d}\subset(I_{S})^{d-n+|S|}$; in particular, it is enough to show that each generator $\ell_{i_{1}}\cdots\ell_{i_{d}}$ of $I_{L,d}$ is in $(I_{S})^{d-n+|S|}$. Since $S$ and $\{i_{1},\ldots,i_{d}\}$ are subsets of $[n]$, we have that $|S\cup\{i_{1},\ldots,i_{d}\}|\leq n$. Also, since $|S|\geq n-d+1$, we have that $n\leq d+|S|-1<d+|S|$. Applying these two facts in succession, we find that:
$|S\cap\{i_{1},\ldots,i_{d}\}|=|S|+d-|S\cup\{i_{1},\ldots,i_{d}\}|\geq|S|+d-n>0.$
But then $S\cap\{i_{1},\ldots,i_{d}\}$ is non-empty and has cardinality at least dn+|S|d-n+|S|. Choose any subset of this intersection that has cardinality dn+|S|d-n+|S|, say $\{j_{1},\ldots,j_{d-n+|S|}\}$. Then each $\ell_{j_{m}}$ is a generator in $I_{S}$ and so the products $\ell_{j_{1}}\cdots\ell_{j_{d-n+|S|}}$ along with $\ell_{i_{1}}\cdots\ell_{i_{d}}$ belong in $(I_{S})^{d-n+|S|}$. ∎
Lemma 2.7.
Let II be the following ideal:
$I=\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}.$
Then $\operatorname{sat}(I)\leq d$.
Proof.
We may assume that $I_{[n]}=\mathfrak{m}$; otherwise, one may work in a polynomial ring with fewer variables. Each term appearing in the decomposition of II is primary. In particular, $(I_{S})^{d-n+|S|}$ is $I_{S}$-primary. Then, a primary decomposition for the saturation $I^{\operatorname{sat}}$ is exactly the same as that of II, except for the removal of ideals that are $\mathfrak{m}$-primary. We have that $\mathfrak{m}^{d}$ appears in the decomposition, since $I_{[n]}=\mathfrak{m}$. Furthermore, the exponent dn+|S|=dn+n=$d$-n+|S|=d-n+n=d is the largest exponent appearing in the decomposition of II. Thus, we see that:
$I=\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}=\mathfrak{m}^{d}\cap\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1,\\ I_{S}\neq\mathfrak{m}\end{subarray}}(I_{S})^{d-n+|S|}=\mathfrak{m}^{d}\cap I^{\operatorname{sat}}.$
This implies that II and $I^{\operatorname{sat}}$ agree in degrees at least $d$, and so $\operatorname{sat}(I)\leq d$. ∎
Theorem 2.8.
We have the following equality of ideals:
$I_{L,d}=\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}.$
Proof.
We proceed by induction on both $d$ and $\dim(R)=r$. We may assume that $K$ is algebraically closed and therefore infinite, as well as that $r$ is minimal so that $I_{[n]}=\mathfrak{m}$.
We first consider the base cases for the induction. Assume that $d=1$. Then $I_{L,d}=\langle\ell_{i}:i\in[n]\rangle=I_{[n]}=\mathfrak{m}$ and:
$\bigcap_{\begin{subarray}{c}S\subset[n],\\ |S|\geq n\end{subarray}}(I_{S})^{1-n+|S|}=I_{[n]}=I_{L,d}.$
Now suppose $d$ is arbitrary and $r=1$. Then $\ell_{i}=a_{i}x_{1}$ with $a_{i}\in K$ not all zero. We then see that $I_{L,d}=\langle x_{1}\rangle^{d}=I^{d}_{[n]}$, and so the claim follows.
We now consider $d,r\geq 2$. In Lemma 2.6, we have shown one of the inclusions, and so the remainder of this proof is dedicated to showing the reverse inclusion. For each $j\in[n]$, define the ideal $I_{j}$ by:
[…]
The induction assumption on $d$ allows us to write the following decomposition for each $I_{j}$:
$I_{j}=\bigcap_{\begin{subarray}{c}S\subset[n]\backslash\{j\},\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}.$
Then:
$\bigcap_{\begin{subarray}{c}|S|\subset[n],\\ |S|\geq n-d+1\end{subarray}}(I_{S})^{d-n+|S|}=\mathfrak{m}^{d}\cap\bigcap_{j=1}^{n}I_{j}.$
To prove the reverse inclusion, it therefore suffices to show that if $f\in\bigcap_{j=1}^{n}I_{j}$ is homogeneous of degree at least $d$, then $f\in I_{L,d}$.
Since $K$ is infinite, there exists a linear form $y\in R$ such that
  • $y$ is a nonzerodivisor on each $I_{j}^{\operatorname{sat}}$ for which $I_{j}$ is positive-dimensional; and
  • $y$ is not in the $K$-span of the linear forms $\ell_{j}$ for all $j$.
Consider the ideals $(I_{L,d}+\langle y\rangle)/\langle y\rangle$ and $(I_{j}+\langle y\rangle)/\langle y\rangle$ in the quotient ring $R/\langle y\rangle$, which has dimension r1r-1. These ideals are generated by $d$ and (d1)(d-1)-fold products of the images of the $\ell_{i}$. One can then invoke the induction hypothesis on $\dim R$ to see that the image of $f$ in $R/\langle y\rangle$ is in $(I_{L,d}+\langle y\rangle)/\langle y\rangle$. Therefore, $f$ can be written as $f=h+yg$, where $h\in I_{L,d}$ and $g\in R$ is of degree at least d1d-1, and so we may assume that $f=yg$. Then, $f=yg\in I_{j}\subset I_{j}^{\operatorname{sat}}$ for each $j$. Since $y$ is a nonzerodivisor on $I_{j}^{\operatorname{sat}}$, we have that $g\in I_{j}^{\operatorname{sat}}$ as well. The preceding argument holds for all $j$ with $I_{j}$ positive-dimensional; if $I_{j}$ has dimension zero, then $I_{j}^{\operatorname{sat}}=R$, and so $g\in I_{j}^{\operatorname{sat}}$ for all $j$. Applying the induction hypothesis along with Lemma 2.7 (so $\operatorname{sat}(I_{j})\leq d-1$), we have that $g\in I_{j}$. Write $y\in\mathfrak{m}=I_{[n]}$ as $y=\sum_{j=1}^{n}y_{j}$, where $y_{j}\in\langle\ell_{j}\rangle$. Then $f=yg=\sum_{j=1}^{n}y_{j}g\in I_{L,d}$, as $y_{j}g\in\langle\ell_{j}\rangle\cdot I_{j}\subset I_{L,d}$. This completes the proof. ∎
@@
We are now able to give a combinatorial description for a primary decomposition of $I_{L,d}$ in terms of the matroid associated with the subspace arrangement, which we now define. The matroid $M$ is on the ground set $n]=\{1,\ldots,n\}$. The independence sets of this matroid are given by subsets $\{i_{1},\ldots,i_{m}\}\subset[n]$ such that the linear forms $\ell_{i_{1}},\ldots,\ell_{i_{m}}$ are linearly independent in the vector space of linear forms in $R$.
@@
A flat of the matroid is defined to be a subset $S\subset[n]$ such that $S=\mathrm{cl}(S)$.
Corollary 2.9.
Let $M$ denote the matroid corresponding to the linear forms $\ell_{1},\ldots,\ell_{n}$. Then we have the following primary decomposition of the ideal $I_{L,d}$:
@@
Proof.
Let $S\subset[n]$ with $|S|\geq n-d+1$. Then $I_{S}=I_{\mathrm{cl}(S)}$, and $|\mathrm{cl}(S)|\geq|S|$, so:
$(I_{S})^{d-n+|S|}\supset(I_{\mathrm{cl}(S)})^{d-n+|\mathrm{cl}(S)|}.$
Therefore, it is enough to only take the closed sets in the intersection of Theorem 2.8; the closed sets are exactly the flats of $M$. ∎
We note that the primary decomposition in Corollary 2.9 is not necessarily minimal, as the following example illustrates. One may try to use more refined data from the associated matroid to determine the associated primes and minimal primary decomposition of $I_{L,d}$, such as in [CT, Theorem 3.2].
@@
Since $S=[n]$ is always a flat, the ideal $\mathfrak{m}^{d}=\langle x,y,z\rangle^{4}$ appears in the primary decomposition given in Corollary 2.9. However, in this example it is redundant and does not appear in the minimal primary decomposition.
@@
which agrees with the primary decomposition given in Corollary 2.9.
We end this subsection with the following result on resolutions and Castelnuovo–Mumford regularity of the ideals $I_{L,d}$. The proof follows the one for [CH, Theorem 3.1] using Lemma 2.7 and Theorem 2.8.
Corollary 2.11.
The ideal $I_{L,d}$ has a linear resolution. In particular, the Castelnuovo–Mumford regularity of $I_{L,d}$ is $d$.
2.3. Affine case
From our work in the homogeneous case, we may now deduce results for primary decompositions of hyperplane arrangements in affine space.
Consider linear polynomials $\ell_{1},\ldots,\ell_{n}\in R$. By homogenizing, we get linear forms $\ell_{1}^{\prime},\ldots,\ell_{n}^{\prime}\in S=K[x_{1},\ldots,x_{r+1}]$. Let $I_{L,d}\subset R$ denote the ideal of $d$-fold products of the $\ell_{i}$, and let $J_{L,d}\subset S$ denote the ideal of $d$-fold products of the $\ell_{i}^{\prime}$. Then:
$I_{L,d}=(J_{L,d}+\langle x_{r+1}-1\rangle)\cap R.$
Therefore, we use Corollary 2.9 to obtain the following primary decomposition of $I_{L,d}$ in $R$:
@@ 2.3. Affine case
Let $\ell_{1},\ldots,\ell_{n}\in R$ be linear polynomials. Then we have the following primary decomposition of $I_{L,d}$:
@@
    the hyperplane at infinity is not in the span of this associated flat, i.e., $x_{r+1}\notin I_{S}$; and
@@
We retain the notation introduced in the paragraph preceding the statement of the theorem. By Corollary 2.9, we have that:
@@
This can be obtained from a primary decomposition for the ideal $J_{L,d}$ of 33-fold products of the $\ell_{i}^{\prime}$, removing ideals supported on the hyperplane $z=0$, and then setting $z=1$:

### v2

@@ Abstract.
We study partial fraction decompositions (PFDs) in several variables using tools from commutative algebra. We give criteria for when a rational function with poles on a hyperplane arrangement has a desirable PFD. Our criteria are obtained by examining the primary decomposition of ideals coming from hyperplane arrangements. We then present an algorithm for finding a PFD that satisfies properties desired for simplifying the calculation of scattering amplitudes. We demonstrate the effectiveness of this algorithm by computing practical examples coming from Feynman integrals.
@@ 1. Introduction
Partial fraction decompositions (PFDs) of rational functions in several variables appear frequently in high-energy physics. Examples include abstract applications, such as the recasting of scattering amplitudes as canonical forms of polytopes [ABL], and computational applications like developing efficient algorithms to compute Feynman integrals [HvM]. Despite their ubiquity in physics, studies of PFDs seldom appear in the mathematical literature [BBD+, Bec, Lei]. There has been interest from physicists to utilize algebro-geometric techniques in order to compute certain PFDs [BW+, CDE, DP]. In this paper we study PFDs from the perspective of commutative algebra, by using primary decomposition and algorithms based upon Gröbner bases. This is done by recasting the existence of PFDs as an ideal membership problem.
@@
In applications, one would like to know when a rational function $f/g$ has a PFD, and if so, how to make the PFD “optimal.” Here, optimal can refer to a number of parameters, including minimizing the number of terms $m$ in the PFD, maximizing the degrees of the numerators $f_{i}$, and avoiding spurious poles (the introduction of such poles can complicate calculations and obscure physical properties and symmetries [Hod]).
@@
Let $f\in K[x_{1},\ldots,x_{r}]$ be a homogeneous polynomial of degree $\deg(f)\geq d\geq 1$, and let $\ell_{1},\ldots,\ell_{n}$ be nonzero linear forms with $n\geq d$. Then $f/(\ell_{1}\cdots\ell_{n})$ has a PFD of degree $d$ as in (1) if and only if the numerator $f$ vanishes to order at least dn+|S|d-n+|S| on all flats $S$ of size at least nd+1n-d+1 of the hyperplane arrangement associated with the linear forms.
Our approach to proving Theorem 1.2 is with commutative algebra: given linear forms $\ell_{1},\ldots,\ell_{n}$ of the polynomial ring $R=K[x_{1},\ldots,x_{r}]$, we use the primary decomposition of the following ideal:
@@
Existence of a PFD of order $d$ for the rational function $f/(\ell_{1}\cdots\ell_{n})$ is equivalent to the numerator $f$ being in the ideal $I_{L,d}$, and a primary decomposition of $I_{L,d}$ gives an equivalent geometric condition for when this occurs.
One can use the computer algebra software Macaulay2 [M2] to determine the existence of and algorithmically compute PFDs of the form in (1). Our algorithm (Algorithm 2) for finding a PFD is based upon Gröbner basis algorithms. It satisfies conditions (i), (ii), (iii), and (iv) of the “wishlist” for a good partial fraction algorithm, described by Heller and von Manteuffel in [HvM, §2.3] and based upon work of Leĭnartas [Lei]:
@@
We now discuss the organization and contributions of this paper. In §2, we discuss the ideals $I_{L,d}$ in (2) defined from hyperplane arrangements. We present primary decompositions for these ideals (Theorems 2.5 and 2.7), which are given in terms of the matroid associated with the hyperplane arrangement. We then apply these decompositions in §3 to study the ideals $I_{L,d}$ where $L$ is the braid arrangement. In §4, we also use the primary decomposition to deduce vanishing behavior of adjoint polynomials arising from polytopes; these polynomials are objects of much study in physics, combinatorics, and algebraic geometry. We then return to our main motivation of PFDs in §5: in this section, we give criteria for when PFDs of degree $d$ exist (Theorems 5.2 and 5.3). We present Algorithms 1 and 2 to determine the existence of and compute PFDs; we also demonstrate how our algorithms satisfy all four conditions (i)(iv) of a good PFD algorithm. In §6, we then show how our criteria and algorithm can be applied to PFDs coming from two different contexts from physics: Feynman integrals (§6.1) and flat-space wavefunction coefficients (§6.2). We conclude in §7 with remarks on the outlook and directions for future investigation for the algebraic study of PFDs.
@@
Remark 1.3.
After initially posting our paper, we were informed that Theorem 2.5 was already proven in [BTX, Toh]. We thank an anonymous referee and S , tefan Tohăneanu for referring us to these results. Our primary contribution, connecting PFDs to commutative algebra, remains original.
@@ 2. Primary decompositions
Let $d$ be a positive integer with $1\leq d\leq n$. Given a hyperplane arrangement $L=(\ell_{1},\ldots,\ell_{n})$, we consider the ideal $I_{L,d}\subset R$ defined by $d$-fold products of the linear forms:
@@
We allow for the cases that some of the $\ell_{i}=0$ are zero, or that the $\ell_{i}$ may not all be distinct, i.e., $\ell_{i}=\ell_{j}$ for for some $i\neq j$. However, we assume that $I_{L,d}\neq 0$ is not the zero ideal, so not all of the $\ell_{i}=0$ are zero. The purpose of this section is to present primary decompositions for $I_{L,d}$, first in the case of a generic hyperplane arrangement and then for arbitrary arrangements.
2.1. Generic hyperplane arrangements in projective space
@@
The main result of this subsection (Theorem 2.3) shows that $I_{L,d}=\mathfrak{m}^{d}$ is the $d$th power of the homogeneous maximal ideal $\mathfrak{m}\subset R$. This was previously proven by Tohăneanu in the context of coding theory [Toh, Theorem 3.1]. We give an alternative proof.
To state the following proposition, we introduce some notation. Write each linear form $\ell_{i}\in R$ in the form $\ell_{i}=\sum_{j=1}^{r}a_{ij}x_{j}$, with $a_{ij}\in K$. Define $A=(a_{ij})$ to be the $(n\times r)$-matrix of these coefficients. For $1\leq i_{1}<\cdots<i_{r}\leq n$, denote by $A[i_{1},\ldots,i_{r}]$ the $r$-minor of $A$ obtained by taking the submatrix of $A$ with rows labeled by $\ell_{i_{1}},\ldots,\ell_{i_{r}}$. Now, define $\mathcal{M}$ to be the $(\binom{n}{d}\times\binom{n}{d})-$matrix of coefficients, with rows labeled by $d$-fold products $\ell_{i_{1}}\cdots\ell_{i_{d}}$ of the linear forms, and columns labeled by degree $d$ monomials of $R$.
@@
The flats of this matroid are the subsets of the ground set of cardinality at most $r$, together with the ground set itself. This foreshadows results of the following subsection, where we discuss arbitrary hyperplane arrangements.
2.2. Arbitrary hyperplane arrangements
In this subsection, we give a combinatorial description for a primary decomposition of $I_{L,d}$. This is given in terms of a matroid associated with the hyperplane arrangement.
@@ 2.2. Arbitrary hyperplane arrangements
We first focus on the case where $L$ is a hyperplane arrangement in projective space. Suppose $L=(\ell_{1},\ldots,\ell_{n})$ are linear forms in $R$. The associated matroid $M$ is on the ground set $n]=\{1,\ldots,n\}$. The independence sets of this matroid are the subsets $\{i_{1},\ldots,i_{m}\}\subset[n]$ such that the linear forms $\ell_{i_{1}},\ldots,\ell_{i_{m}}$ are linearly independent in the vector space of linear forms in $R$.
@@
A flat of the matroid $M$ is defined to be a subset $S\subset[n]$ such that $S=\mathrm{cl}(S)$.
The primary decompositions we consider are defined in terms of powers of the prime ideals
$I_{S}=\langle\ell_{i}:i\in S\rangle,\qquad S\subset[n].$
These primary decompositions are given in [BTX, Theorem 2.2]. The following is a reformulation of their result.
Theorem 2.5.
Let $M$ denote the matroid corresponding to the linear forms $L=(\ell_{1},\ldots,\ell_{n})$. Then we have the following primary decomposition of the ideal $I_{L,d}$:
@@
We note that the primary decomposition in Theorem 2.5 is not necessarily minimal, as the following example illustrates. One may try to use more refined data from the associated matroid to determine the associated primes and minimal primary decomposition of $I_{L,d}$, such as in [CT, Theorem 3.2].
@@
Since $S=[n]$ is always a flat, the ideal $\mathfrak{m}^{d}=\langle x,y,z\rangle^{4}$ appears in the primary decomposition given in Theorem 2.5. However, in this example it is redundant and does not appear in the minimal primary decomposition.
@@
which agrees with the primary decomposition given in Theorem 2.5.
From the homogeneous case, we may now deduce results for primary decompositions of hyperplane arrangements in affine space.
Consider linear polynomials $\ell_{1},\ldots,\ell_{n}\in R$. By homogenizing, we get linear forms $\ell_{1}^{\prime},\ldots,\ell_{n}^{\prime}\in R^{\prime}=K[x_{1},\ldots,x_{r+1}]$. Let $I_{L,d}\subset R$ denote the ideal of $d$-fold products of the $\ell_{i}$, and let $J_{L,d}\subset R^{\prime}$ denote the ideal of $d$-fold products of the $\ell_{i}^{\prime}$. Then:
$I_{L,d}=(J_{L,d}+\langle x_{r+1}-1\rangle)\cap R,$
where $J_{L,d}$, via Theorem 2.5, has a primary decomposition in terms of powers of ideals $J_{S}=\langle\ell_{i}^{\prime}:i\in S\rangle\subset R^{\prime}$. This enables us to obtain the following primary decomposition of $I_{L,d}$ in $R$:
@@
Let $L=(\ell_{1},\ldots,\ell_{n})$ with $\ell_{i}\in R$ be linear polynomials. Then we have the following primary decomposition of $I_{L,d}$:
@@
    the hyperplane at infinity is not in the span of this associated flat, i.e., $x_{r+1}\notin J_{S}$; and
@@
We retain the notation introduced in the paragraph preceding the statement of the theorem. By Theorem 2.5, we have that:
@@
This can be obtained from a primary decomposition for the ideal $J_{L,d}$ of $3$-fold products of the $\ell_{i}^{\prime}$, removing ideals supported on the hyperplane $z=0$, and then setting $z=1$: